OpenStudy (anonymous):

Johnson Canniver has a repair shop. He bought a used lawn mower for $35.00 from a customer. He spent $39.17 on parts to repair the mower. He also painted and polished the mower at a cost of $36.43. He was able to sell the refurbished lawn mower for $155.00. If it took 4.5 hours to do all the work, how much did he earn per hour to the nearest cent?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

I got 9.98 but its wrong

OpenStudy (johnweldon1993):

Okay I feel like this is a trick question...are we supposed to include the 35 it cost for him to buy it? because that doesn't REALLY count as working on it...hmm

OpenStudy (anonymous):

lets try it without it?

OpenStudy (johnweldon1993):

Alright lol So lets see....total work cost would be the 39.17 + 36.43 = 75.6 for 4.5 hours to do all that we have 75.6 / 4.5 = 16.8 per hour that he got "paid" This is JUST for the work

OpenStudy (johnweldon1993):

However if we consider this ENTIRE process to be included...we have the cost of work + cost of original = 110.6 And he also sold it for 155 so that would still be the 44.4 in profit If all he "earned" from doing all the work was 44.4 dollars...in 4.5 hours that would be 9.867 dollars per hour

OpenStudy (anonymous):

it says it wrong hmm

OpenStudy (johnweldon1993):

Both are wrong?

OpenStudy (anonymous):

2) Need to add up the total cost. 35 (the purchase price) + 39.17 (spent on parts) + 36.43 (paint and polish) = $110.06 the amount of $ JC spent on the mower. 110.06/4.5 = $24.46 per hour SPENT on mower. Sold it for 155. 155/4.5 = $34.44 per hour earned on mower. 34.44 - 24.46 = $9.98 per hour earned on mower. yes this is another answer i got but its wrong

OpenStudy (anonymous):

im typing it on the compute online class

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i liked john idea, buying it increases its cost, but not its labor perse let me chk your mathings

OpenStudy (amistre64):

35 to buy money spent on purchase not on labor 39.17 on parts to repair the mower money spent on parts doesnt tell us about labor according to johns idea 36.43 for painted and polished the mower. labor 155.00 sold; it took 4.5 hours

OpenStudy (amistre64):

(155-36.43)/4.5 26.35 ?

OpenStudy (anonymous):

nop

OpenStudy (amistre64):

im not sure what they want to be honest. this seems a bit course related like if it was in management or accounting

OpenStudy (anonymous):

business math

OpenStudy (amistre64):

your material should have a similar example.

OpenStudy (johnweldon1993):

^That's why I tried looking for examples...but got nothing...sorry @fani1996 :(

OpenStudy (anonymous):

its ok :) thanks guys

OpenStudy (amistre64):

155/4.5 is my only other thought but then this is just guessing at this point ....